sum of three consecutive numbers is eight more than twice the lowest number. What are the three numbers
Answer by amarjeeth123(569)

Let the consecutive numbers be x-1,x,x+1.
sum of three consecutive numbers is eight more than twice the lowest number.
x-1+x+x+1=8+2(x-1)
3x=8+2x-2
x=6
The three numbers are 5,6 and 7.
Verifying the solution:
5+6+7=18
8+2(5)=18.